# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ 4.3 Key Findings
Given the findings of this SLR, we attempt to address our Research Questions.
@@
Phishing URLs. While established datasets have played a crucial role in the development of phishing URL detection models, there is a clear need to incorporate more dynamic and current data sources. Leveraging user-reported phishing URLs from social networks, along with data from Telecom and Security organizations, offers a more effective approach to combating phishing attacks. These sources provide real-time, diverse, and relevant data that enhance the robustness and accuracy of detection models, keeping pace with the evolving nature of phishing threats. By combining the strengths of both traditional and modern data sources, researchers can develop more comprehensive and adaptive phishing detection systems, better protecting users from phishing URLs.
@@
Regarding the methodologies used, we find that the existing literature used state-of-the-art methodologies for the analysis and detection of phishing URLs. Notably, RF seems to be the stand-alone model that works best, while other hybrid methodologies also report promising performance.
@@
Phishing Emails. While publicly available datasets have laid the groundwork for phishing email detection research, the rapidly evolving nature of phishing attacks requires the use of more dynamic and up-to-date data sources. Leveraging user-reported emails, real-time spam collections, and advanced synthetic data generation techniques can significantly enhance the robustness and accuracy of phishing detection models. By combining traditional datasets with innovative data sources, researchers can develop more comprehensive and adaptive phishing detection systems, better equipped to detect phishing activities via email.
@@
All of the works that opted for automated phishing email detection report very good performance on their detection models, with RF, BERT, LSTM, RNN, and SVM being the most popular.
@@
Phishing SMS By this point we established that many works in various scam type analysis rely on established datasets. Alas, the rapidly evolving nature of smishing requires a more dynamic and diversified approach to data collection. By integrating publicly available datasets with real-time, user-reported data and specialized security sources, researchers can develop more effective and resilient smishing detection models. This approach will ensure that models remain relevant and capable of addressing new and sophisticated smishing threats as they arise.
@@
Contrary to the phishing email detection, we find that in the case of phishing SMS detection, which tends to be much smaller text, SVM and various applications of Neural Networks perform best.
@@
Phishing Phone Calls. Regarding the data sources used for automated vishing detection, most works used text data via transcribing voice recordings. Other works also used caller ID information from various telecommunication operators. Some works collected data from user reports and only one work attempted to detect deepfake signals in voice recordings. The best performing methodology used for automated vishing detection is SVM.
@@
Phishing User Reports Reviewing the four works that studied phishing via user reports, we find that researchers used data from various sources, including court judgements, public data from forums, and user reports from Financial Institutions. The methodologies applied vary as there is use of Named Entity Recognition, various NLP methodologies, and AI methodologies.
@@
Fake Reviews. Although many works used previously available datasets to establish and compare their detection models, we notice a clear trend where later studies tend to collect data from platforms like Amazon, Google Play, Apple App Store, and YouTube for their analysis. This is extremely encouraging as the data used for these detection models need to be constantly updated. Hybrid models including LR, SVM, CNN, and LSTM seem to perform best for the detection of fake reviews.
@@
Fraudulent Job Postings. While the Kaggle dataset has been pivotal in advancing research on fraudulent job postings, the rapidly evolving nature of job scams necessitates the use of more current and diverse data sources. Custom data collection methods, which tap into active job posting sites, represent a crucial step forward in enhancing the effectiveness of detection models. By leveraging a mix of established and new data sources, researchers can develop more comprehensive systems to effectively combat fraudulent job postings. The models that perform best for this fraudulent activity vary. We find that Bi-LSTM, KNN, RF, DNN, and LightGBM perform well.
@@
GenAI Social Engineering. Only two works here use data collected from real use cases, namely Ayoobi et al. ayoobi2023looming discuss fake profile AI-generated content on professional social networks, while Diresta et al. diresta2024spammers present deep fakes posted on Facebook. Other works discuss the potential sybersecurity, privacy, and ethical issues with these models and how they can be misused to automatically create fraudulent content.
@@ 5 Discussion
We now discuss our findings, detailing recognised limitations and shortcomings detected in the reporting of AI models related to the performance and data sources used. We also provide recommendations for researchers developing detection models for online fraud.
@@ 5.1 Data Sources
Overall, we analyzed the data sources and detection methodologies of 222 papers that aim to address a range of online fraud problems. Although our findings reveal a preference for well-established datasets, especially in the automated detection of various phishing and fake reviews detection, more recent studies (published after 2023) seem to shift towards more dynamic and recent sources.
@@
We found that the overwhelming majority of the Phishing domain detection studies relied on well-known and extensively studied datasets from websites like PhishTank, OpenPhish, and SpamHaus. In contrast, others used datasets made available from previous studies or publicly available repositories like Kaggle, University of California Irvine (UCI) Machine Learning Repository, and GitHub. This is also the case for studies that focused on phishing email detection, fake review detection, and fraudulent recruitment detection. However, while these established datasets provide a valuable foundation for research, they come with limitations.
@@
Online fraud is dynamic, with new scam techniques frequently emerging and older methods, like phishing URLs and websites, continually evolving. Studies show hat LLM empowered bots or scammers could be deployed to generate and automate sophisticated and targeted fraudulent and phishing content online, either that being an email, a professional profile, or deceptive Terms and Conditions for fake e-commerce websites. Hence, relying on outdated datasets may limit the effectiveness of detection models when applied to current or evolving threats. At the same time, the static nature of previously used datasets does not capture the latest trends and variations of various online fraud techniques and activities.
@@
To address these limitations, recent studies have started leveraging more dynamic and real-time data sources. Regarding automated phishing detection, recent studies used user-reported phishing URLs from social networks, as well as data from Telecom and Security organizations. For instance, studies published in 2023 utilized data from sources like Twitter, Facebook, and specialized security institutions bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real ; mehndiratta2023malicious ; yu2024efficient ; liang2019using .
@@
At the same time, recent works on automated phishing email detection have utilised user-reported emails, providing a real-time perspective on phishing threats gallo2019identifying . Genc and Jianggenc2021understanding ; jiang2024detecting used emails from their personal or professional spam folders, capturing a more realistic and up-to-date snapshot of phishing attacks. Mehdi et al.mehdi2023adversarial took an innovative approach by incorporating various techniques to develop their dataset; GPT-2 generated synthetic phishing emails, along with tools like TextAttack, TextFooler, and PWWS. This approach not only provides a diverse dataset but also ensures that the model is robust against sophisticated phishing techniques. Janez et al. janez2021trustworthiness used data from the SPAM Archive [^28] http://untroubled.org/spam/, a continuously updated repository of spam emails.
@@
Similarly, several studies on automated phishing SMS detection have extended their datasets by combining pre-existing publicly available datasets with additional sources to improve the robustness and generalizability of their models, like incorporating additional data from Fake Base Stations zhang2020lies , emails and YouTube comments vinothkumar2022detection , user reported data for research lai2022semorph or on Twitter tang2022clues . Notably, Timko et al. timko2023commercial proposed a platform, SmishTank, where users can post phishing SMS messages, creating an ongoing and up-to-date repository for researchers. We also observed studies using data from specialized security agencies seo2024device and mobile security services liu2021detecting , which offer a more targeted collection of smishing examples.
@@
The overwhelming amount of studies focus on Fraudulent Recruitment detection using the same dataset. In contrast, three studies employed custom crawlers to gather data from various job posting websites in different countries, providing a more diverse and up-to-date perspective. Mahbub et al. mahbub2022online collected data from job posting sites in the UK, Tabassum et al. tabassum2021detecting gathered job postings from Bangladeshi sites, and Zhang et al. zhang2023orfpprediction sourced data from Chinese job sites. The use of up-to-date data collection from these sources offer some advantages. For one, by scraping data from active job posting sites, these studies can capture the most recent and relevant data, reflecting current fraudulent practices. Secondly, collecting data from multiple sources across different regions provides a richer and more varied dataset, which can enhance the robustness and generalizability of detection models. Finally, custom datasets often include a wider variety of job postings, including niche or less common types of employment scams, which can be critical for developing more comprehensive detection systems.
@@
  • Real-time Updates: Social networks and security organizations provide data that is continuously updated. This ensures that the detection models are trained on the most recent phishing URLs, making them more robust against new and emerging threats.
@@
However, despite the advancements in data collection methods, there are still gaps. For instance, the data sources for various works on various online fraud categories were not clearly stated within the manuscripts. This lack of transparency can hinder reproducibility and the ability to compare results across different studies.
@@ 5.2 Methodologies
The methodologies employed across various studies on phishing and fraudulent activities involve a wide range of AI that incorporate NLP, machine learning and deep learning techniques. Most of these techniques involve using extracting features using NLP techniques and then applying supervised machine learning (i.e. labelled data) and deep learning algorithms to build binary classifiers.
@@
For phishing URLs, Machine Learning and Deep learning algorithms such as CNN, ANN, KNN, LSTM, NB, RF, DT, SVM, and XGBoost were commonly used, often with URL feature extraction through NLP methods like character counts and n-grams. We also found various works that applied hybrid approaches where they combined multiple methodologies, demonstrating strong detection capabilities. Similarly, phishing email detection heavily relied on NLP for feature extraction (e.g., LDA, BERT) followed by AI models like RF, NB, and SVM for classification. Similar techniques were applied for phishing SMS detection.
@@
In vishing (phishing phone calls), transcript analysis primarily conducted using NLP and AI models, with some studies examining deepfake voice detection. User reports on phishing activities utilized models like BERT and RF, while fake reviews often involved sentiment analysis with methods like VADER.
@@
Fraudulent job postings detection involved both use of machine learning and deeep learning models such as LR and Bi-LSTM. Romance fraud detection used sentiment detection methods and machine learning models (e.g. RF) and deep learning models (e.g. LSTM). For fraudulent investment and crypto manipulation, classic machine learning models like DT, XGBoost, and SVM were employed. Studies on fraudulent e-commerce and crowdfunding leveraged advanced NLP and machine learning techniques, including GPT-4 and LR, respectively.
@@
Although the research and detection methodologies applied by the detected literature perform well, they are not without theirlimitations. Overall, popular machine learning algorithms like RF and SVM often rely heavily on the quality of features extracted, which can be labor-intensive and may miss out on subtle indicators when dealing with large set of data.
@@
In addition, complex models based on deep learning techniques like Bi-LSTM with VGG or hybrid approaches can be computationally expensive and difficult to implement in real-time systems due to the large amount of data processing resources required.
@@
Notably, models developed that work well for one kind of fraud, might not generalize well on different fraud activities. For example, SMS messages are typically short, providing limited data for accurate feature extraction and classification. In addition, natural language processing techniques used may struggle to capture the features for extracting the understanding the context (semantics) or syntax of phishing content or , leading to potential false positives/negatives (i.e. incorrect misclassificaitons).
@@
Models trained on specific languages or datasets may not perform well on emails in other languages or different styles. Many of the challenges that may arise regarding the trained AI models, are often result of the poor quality of data. More specifically, effective feature extraction is critical but can be difficult due to the varied nature of how natural language is used. At the same time, the textual content used in online fraud activities - such as fraudulent email, sms, or job posting - keeps evolving, making it challenging for static models to remain effective over time.
@@
Although collecting data from various websites, forums, social networks, and telecommunication operators for online fraud detection is invaluable, at the same time, different platform hold data inconsistently, or have unique features and user behaviors, complicating model generalization. Also, identifying fraudulent activities in real-time is challenging due to the dynamic nature of these scams and the lack of real-time observance in the various data sources.
@@
Overall, while these methodologies offer powerful tools for detecting phishing and fraudulent activities, they have challenges related to feature extraction, model complexity, generalizability, and data quality. Finally, one of the major issues higlighted in many of these studies is that these models are using supervised machine learning models that require labeled data. Creating labeled data is often challenging and time-consuming. As a result, researchers have been using existing labeled data, which may become less effective as fraud is evolving over time (e.g. content of phishing emails changing).

### v2

@@ 5.3 Key Findings
We now summarise our findings in relation to the research questions listed in Section 2. To remind the reader, these questions were: to detect the state-of-the-art AI techniques used to detect online fraud (RQ1); the data sources used (RQ2); how researchers evaluate their AI models (RQ3), and what the most studied fraud activities were (RQ4). The answers to RQs 1-3 are organised by fraud type, while the listing of these fraud crime types addresses RQ4.
@@
Takeaways: Phishing URLs. While established datasets have played a major role in developing phishing URL detection models, there is a clear need to incorporate more dynamic and current data sources. Leveraging user-reported phishing URLs from social networks and data from telecom and security organizations would offer a more effective approach to combating phishing attacks. These sources provide real-time, diverse, and relevant data that enhance the robustness and accuracy of detection models, keeping pace with the evolving nature of phishing threats. By combining the strengths of both traditional and modern data sources, researchers can develop more comprehensive and adaptive phishing detection systems, better protecting users from phishing URLs.
@@
Regarding the methodologies used, we find that the existing literature used state-of-the-art methodologies to analyse and detect phishing URLs. Of these, RF seems to be the stand-alone model that works best, while other hybrid methodologies also report promising performance. Regarding performance reporting, authors often fail to adequately report all of the performance metrics of their model. Although the Accuracy of the model is reported in all but seven studies, other metrics like Precision, Recall, F1, and AUC are omitted in more than half of the studies we analysed for this type of online fraud.
@@
Takeaways: Phishing Emails. While publicly available datasets have laid the groundwork for phishing email detection research, the rapidly evolving nature of phishing attacks requires the use of more dynamic and up-to-date data sources. Leveraging user-reported emails, real-time spam collections, and advanced synthetic data generation techniques could significantly enhance the robustness and accuracy of phishing detection models. By combining traditional datasets with innovative data sources, researchers could develop more comprehensive and adaptive phishing detection systems that are better equipped to detect phishing activities via email.
@@
All of the studies that developed automated approaches to phishing email detection reported very good performance, with RF, BERT, LSTM, RNN, and SVM being the most popular. Similar to the Phishing URLs above, accuracy was the metric reported most often. Only two studies reported AUC, and Precision, Recall, and F1 were rarely reported.
@@
Takeaways: Phishing SMS. This type of scam was also found to rely on existing datasets. Alas, the rapidly evolving nature of smishing requires a more dynamic and diversified approach to data collection. Researchers could develop more effective and resilient smishing detection models by integrating publicly available datasets with real-time, user-reported data and specialized security sources. This approach would ensure that models remain relevant and capable of addressing new and sophisticated smishing threats as they arise.
@@
Contrary to phishing email detection, we find that in the case of phishing SMS detection, which tends to involve much less text, SVM and various applications of Neural Networks performed best. Again, the Precision, Recall, F1-score, and AUC metrics were underreported, with Accuracy being the metric most studies report.
@@
Takeaways: Phishing Phone Calls. Regarding the data sources used for automated vishing detection, most studies used text data obtained by transcribing voice recordings. Other research also used caller ID information from various telecommunication operators. Some researchers collected data from user reports, and only one attempted to detect deepfake signals in voice recordings. The best-performing technique used for automated vishing detection was SVM. Although Accuracy was also the most reported performance metric for this category, many studies failed to provide any metrics.
@@
Takeaways: Phishing (User Reports). Reviewing the four studies that focused on phishing via user reports, we found that the researchers used data from various sources, including court judgements, public data from forums, and user reports from Financial Institutions. The applied methodologies varied and included Named Entity Recognition, various NLP and statistical techniques (D2V, Jaccard, TF-IDF), and ML techniques (SMO, J48, NB, RF, XGBoost). Only one study liu2021fine provided performance metrics, and this was for their BERT model that performed best.
@@
Takeaways: Fake Reviews. Although many studies used previously available datasets to establish and test their detection models, we noticed a clear trend where more recent studies tended to collect data from platforms like Amazon, Google Play, the Apple App Store, and YouTube. This is very encouraging as the data used for these detection models need to be constantly updated. Hybrid models, including LR, SVM, CNN, and LSTM, seemed to perform best in the detection of fake reviews. Again, the Precision, Recall, F1, and AUC metrics were underreported, with Accuracy being the metric most studies report.
@@
Takeaways: Recruitment Fraud. While the Kaggle dataset has been pivotal in advancing research on fraudulent job postings, the rapidly evolving nature of job scams necessitates the use of more current and diverse data sources. Custom data collection methods, which tap into active job posting sites, represent a critical step forward in enhancing the effectiveness of detection models. Researchers can develop more comprehensive systems to effectively combat fraudulent job postings by leveraging a mix of established and new data sources. The models that performed best for this fraudulent activity varied. However, we find that Bi-LSTM, KNN, RF, DNN, and LightGBM performed well. All but one study reported the accuracy performance metric of their best-performing model, while the other four metrics (Precision, Recall, F1, and AUC) remain underreported.
Takeaways: Fake Accounts. Most studies utilised user profile data from popular social networks such as Twitter, Instagram, Facebook, YouTube, and Sina Weibo. Several works relied on openly accessible datasets published from previous studies, and others collect user data through APIs, web scraping, or manual curation. Advanced techniques such as RoBERTa, CatBoost, and Graph Convolutional Neural Networks (GCNN) have been employed alongside traditional classifiers like Random Forest, Support Vector Machines, and Neural Networks, with Random Forest being one of the most popular and frequently high-performing models.
Performance across studies is generally strong, with reported accuracy often exceeding 90%, though other metrics like Precision, Recall, and F1 are underreported in some works. However, the rapidly changing strategies used by fake account creators highlight the need for more dynamic datasets and adaptive models to tackle this challenge effectively.
@@
Takeaways: GenAI Social Engineering Attacks. Many studies discuss how GenAI models might affect cybersecurity and privacy, the ethical issues they pose, and how they could be misused to create fraudulent content automatically. However, only two studies identified in our review used data collected from real use cases. These were Ayoobi et al. ayoobi2023looming , who discussed fake profile AI-generated content on professional social networks, and Diresta et al. diresta2024spammers , who examined deepfakes posted on Facebook. No performance metrics were reported in these studies as they were not applicable. Authors studying this offence were not building models but rather evaluating or experimenting with existing tools, including OpenAI’s ChatGPT.
Takeaways: Social Media Scams. The detection of social media scams presents unique challenges due to the diversity of platforms, fraud types, and datasets. Most studies leverage platform-specific data sources such as Telegram la2023sa ; la2021uncovering ; shah2020illicit , WeChat xu2022efficiently , Twitter al2023dspamonto ; shah2020illicit , and YouTube tripathi2022analyzing to build detection models.
The methodologies employed in these studies vary widely, encompassing advanced NLP techniques (W2V, P2V, and LDA) shah2020illicit ; la2021uncovering , as well as machine learning classifiers (RF, NB, and SVM) tripathi2022analyzing ; al2023dspamonto .
While these studies report promising results, this category involved studies on various kinds of social media scams. Hence, the lack of standardisation across models and datasets limits the generalisability of these findings. Notably, these studies also often underreported evaluation metrics like AUC and F1.
Takeaways: Romance Fraud. Romance fraud detection has primarily focused on analysing user-generated content on dating platforms and social media. Studies utilised diverse datasets, including user profiles from dating platforms he2021datingsec ; suarez2019automatically and tweets tagged with #tinderswindler lokanan2023tinder . Methods employed feature extraction techniques like BOW, TF-IDF, and sentiment analysis with textBlob lokanan2023tinder . Among machine learning models, RF consistently performed well lokanan2023tinder , while LSTM achieved the best performance in detecting malicious accounts in dating apps he2021datingsec , and EML yielded high accuracy in identifying fraudulent profiles suarez2019automatically . Due to the limited number of studies under this category, we cannot make conclusions regarding the models’ overall performance reporting and generalisability.
Takeaways: GenAI Scam Baiting. The use of GenAI for scam baiting has shown promising potential in wasting scammers’ resources and collecting data for fraud analysis. Studies employed LLMs such as ChatGPT cambiaso2023scamming ; bajaj2023automatic and semi-unsupervised DistillBERT chen2023active ; bajaj2023automatic to engage with scammers and categorise phishing emails. These studies highlighted the potential of GenAI tools in automating scam baiting, but future work should focus on creating publicly available datasets and refining engagement strategies to improve scalability and efficacy.
Takeaways: Fraudulent Investment Detection. The detection of fraudulent investment scams has been explored using a variety of data sources, including Bitcointalk siu2022invest , YouTube li2023towards , and instant messaging apps kuo2024constructing . However, the small sample of just three studies limits the ability to conclusively identify the best-performing model for this category. The studies reported varied performance metrics, with Siu et al. siu2022invest highlighting XGBoost as the top performer for detecting fraudulent investment advertisements on Bitcointalk. Kuo and Tsang kuo2024constructing found DT to be the best model for identifying emotional fluctuations in scam discussions on a popular Taiwanese messaging app, and Li et al. li2023towards focused on arbitrage bot scams and utilised NNs for their analysis, without reporting performance metrics. As the three studies under this category analysed different aspects of a fraudulent investment, we cannot draw any conclusions regarding the best-performing model.
Takeaways: Fraudulent Crowdfunding. Fraudulent crowdfunding detection has been explored using a small set of two studies, both focusing on Kickstarter campaigns lee2022backers ; shafqat2019topic . LR, in combination with many NLP features, was employed to identify fraudulent campaigns, achieving an accuracy of 87.3%. A combination of LSTM and LDA based on user comments was also reported, without reporting performance metrics. Given the limited scope of these studies, further research is required to assess the robustness of these models in detecting fraudulent crowdfunding activities. Both studies highlight the effectiveness of NLP-based approaches for feature extraction and classification in this domain.
Takeaways: Crypto Manipulation. The detection of cryptocurrency market manipulation was investigated in two studies, which used data collected from Twitter, Telegram, and Discord nizzoli2020charting ; mirtaheri2021identifying . Methodologies used included network analysis, topic analysis, and SVM. The limited number of studies made it difficult to assess the robustness of the models, and further research is needed to identify the best-performing AI technique for this type of fraud.
Takeaways: Fraudulent E-Commerce. The only study identified for fraudulent e-commerce detection salihovic2019role used OpenAI’s GPT-4 model to analyze the terms and conditions of e-commerce websites and detect obscure financial obligations, such as shipping, subscription, and refund fees. However, the study did not provide performance metrics or compare different models. Due to the limited nature of this single study, it is impossible to draw conclusions about the approach’s effectiveness or the model’s general applicability in detecting fraudulent e-commerce websites.
@@ 6 Discussion
We now discuss our findings, discussing recognised limitations and shortcomings identified in the reporting of AI models related to the performance and data sources used. We also provide recommendations for researchers developing detection models for online fraud.
@@ 6.1 Data Sources
Overall, we analysed the data sources used and the detection methodologies employed in $223$ papers that aimed to address a range of online fraud problems. Although our findings reveal a preference for well-established datasets, especially in the automated detection of various phishing and fake reviews detection, more recent studies (published after 2023) seem to shift towards more dynamic and recent sources.
@@
We found that the overwhelming majority of studies concerned with the detection of phishing relied on well-known and extensively studied datasets from websites like PhishTank, OpenPhish, and SpamHaus. In contrast, others used datasets made available from previous studies or publicly available repositories like Kaggle, the University of California Irvine (UCI) Machine Learning Repository, and GitHub. This was also the case for studies that focused on phishing email detection, fake review detection, and fraudulent recruitment detection. While these established datasets provide a valuable foundation for research, they come with limitations.
@@
Online fraud is dynamic, with new scam techniques continually evolving, or building over older scam methods. Studies show that LLM-empowered bots, or scammers could be deployed to generate and automate sophisticated and targeted fraudulent and phishing content online, either in the form of email, a professional profile, or deceptive terms and conditions for fake e-commerce websites janjeva2023rapid ; schmitt2023digital ; ferrara2024genai . Hence, relying on outdated datasets may limit the effectiveness of detection models when applied to current or evolving threats. The historical datasets used do not capture the latest trends and variations in the different online fraud techniques and activities we see today (and will see tomorrow).
@@
Recent studies have started to leverage more dynamic and real-time data sources to address these limitations. Regarding automated phishing detection, recent studies used user-reported phishing URLs from social networks and data from telecom and security organizations. For instance, studies published in 2023 utilized data from sources like Twitter, Facebook, and specialized security organisations bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real ; mehndiratta2023malicious ; yu2024efficient ; liang2019using .
@@
At the same time, recent research on automated phishing email detection has utilised user-reported emails, providing a real-time perspective on phishing threats gallo2019identifying . For example, Genc and Jianggenc2021understanding ; jiang2024detecting used emails from their personal or professional spam folders, capturing a more realistic and up-to-date snapshot of phishing attacks. Mehdi et al.mehdi2023adversarial took an innovative approach by incorporating various techniques to develop their dataset; GPT-2 generated synthetic phishing emails and tools like TextAttack, TextFooler and PWWS. This approach provides a diverse dataset and ensures the model is robust against sophisticated phishing techniques. In their study, Janez et al. janez2021trustworthiness used data from the SPAM Archive [^31] http://untroubled.org/spam/, which is a continuously updated repository of spam emails.
@@
Similarly, several studies concerned with automated phishing SMS detection have extended the datasets used by combining pre-existing publicly available datasets with additional sources to improve the robustness and generalizability of their models. This has included the use of additional data from Fake Base Stations zhang2020lies , emails, YouTube comments vinothkumar2022detection , user-reported data for research lai2022semorph or content posted on Twitter tang2022clues . In particular, Timko et al. timko2023commercial proposed a platform, SmishTank, where users can post phishing SMS messages, creating an ongoing and up-to-date repository for researchers. We also observed studies using data from specialised security agencies seo2024device and mobile security services liu2021detecting , which offer a more targeted collection of smishing examples.
@@
An overwhelming number of studies focused on fraudulent recruitment detection using the same dataset from Kaggle. [^32] https://www.kaggle.com/datasets/amruthjithrajvr/recruitment-scam Only three studies employed custom crawlers to gather data from various job posting websites in different countries, providing a more diverse and up-to-date perspective. Mahbub et al. mahbub2022online collected data from job posting sites in the UK, Tabassum et al. tabassum2021detecting gathered job postings from Bangladeshi sites, and Zhang et al. zhang2023orfpprediction sourced data from Chinese job sites. The use of up-to-date data collection from these sources offers some advantages. For one, these studies can capture the most recent and relevant data by scraping data from active job posting sites, reflecting current fraudulent practices. Second, collecting data from multiple sources across different regions provides a richer and more varied dataset, which can enhance the robustness and generalizability of detection models. Finally, custom datasets often include a wider variety of job postings, including niche or less common types of employment scams, which can be critical for developing more comprehensive detection systems.
@@
  • Real-time Updates: Social networks and security organizations provide continuously updated data. This ensures that the detection models are trained on the most recent phishing URLs, making them more robust against new and emerging threats.
@@
However, despite the advancements in data collection methods, there are still gaps. For instance, the data sources used in studies for several types of online fraud were not clearly stated within the manuscripts. This lack of transparency can hinder reproducibility and the ability to compare results across different studies.
@@ 6.2 Methodologies
The methodologies employed across various studies of phishing and fraudulent activities involved a wide range of AI that incorporate NLP, machine learning and deep learning techniques. Most of these techniques involved the extraction of features using NLP techniques and then applying supervised machine learning (i.e. models that use labelled data) and deep learning algorithms to build binary classifiers.
@@
For phishing URLs, Machine Learning and Deep learning algorithms such as CNN, ANN, KNN, LSTM, NB, RF, DT, SVM, and XGBoost were commonly used, often with URL feature extraction through NLP methods like character counts and n-grams. We also found various studies that applied hybrid approaches combining multiple techniques, demonstrating strong detection capabilities. Similarly, phishing email detection heavily relied on NLP for feature extraction (e.g., LDA, BERT) followed by AI models like RF, NB, and SVM for classification. Similar techniques were applied for phishing SMS detection.
@@
For vishing (phishing phone calls), transcript analysis was primarily conducted using NLP and AI models, with some studies examining deepfake voice detection. The analysis of user reports about phishing activities utilized models like BERT and RF, while studies of fake reviews often involved sentiment analysis using methods like VADER.
@@
Fraudulent job posting detection has involved both the use of machine learning and deep learning models such as LR and Bi-LSTM, while romance fraud detection has used sentiment detection methods in combination with machine learning models (e.g. RF) and deep learning models (e.g. LSTM). Classic machine learning models like DT, XGBoost, and SVM were employed for fraudulent investment and crypto manipulation. Studies on fraudulent e-commerce and crowdfunding leveraged advanced NLP and machine learning techniques, including GPT-4 and LR, respectively.
@@
Although the research and detection methodologies applied in the reviewed literature performed well, they are not without limitations. Overall, popular machine learning algorithms like RF and SVM often rely heavily on the quality of extracted features, which can be labour-intensive to generate and may miss out on subtle indicators when dealing with large data sets.
@@
In addition, complex models based on deep learning techniques like Bi-LSTM with VGG or hybrid approaches can be computationally expensive and difficult to implement in real-time systems due to the large amount of data processing resources that they require.
@@
Notably, models developed that work well for one kind of fraud might not be generalised well to other fraud activities. For example, unlike emails, SMS messages are typically short, providing limited data for accurate feature extraction and classification. In addition, natural language processing techniques used may struggle to capture those features necessary to understand the context (semantics) or syntax of phishing content, leading to potential false positives/negatives (i.e., misclassifications).
@@
Models trained on specific languages or datasets may not perform well on emails in other languages or different styles. Many of the challenges that may arise regarding the trained AI models are often the result of poor data quality. More specifically, effective feature extraction is critical but can be difficult due to the varied nature of how natural language is used. At the same time, the textual content used in online fraud activities - such as fraudulent emails, SMSs, or job postings - keeps evolving, making it challenging for static models to remain effective over time.
@@
Although collecting data from various websites, forums, social networks, and telecommunication operators for online fraud detection is invaluable, at the same time, different platforms hold data inconsistently or have unique features and user behaviours, complicating model generalization. Also, identifying fraudulent activities in real-time is challenging due to the dynamic nature of these scams and the lack of real-time occurrences in the various data sources.
@@
In short, while these methodologies offer powerful tools for detecting phishing and fraudulent activities, there are challenges related to feature extraction, model complexity, generalizability, and data quality. Finally, one of the major issues highlighted in many of these studies was that the models use supervised machine learning models, which require labelled data. Creating labelled data is often challenging and time-consuming. This is one reason why so many researchers use existing labelled data, but, as discussed, such data will become less useful as fraud evolves over time.